1.The nurse is caring for a client with a prescription for enoxaparin 2mg/kg subcutaneous daily. The client weighs 155lbs 8oz the pharmacyndispenses enoxaparin 40mg/0.4ml how many ml will the nurse administer. Record answer rounding to the nearest tenth.

2. The nurse is caring for a child with the following prescription dicloxacillin sodium 50mg po every 6hrs the child weighs 36lbs 4oz. The recommended dosage for dicloxacillin sodium for oral suspension is 12-13mg/kg/day for children less than 40mg in equally divided doses every 6hr

a.Is the prescribed amount a safe dose yes or no explain

b. The pharmacy has dispensed dicloxacillin sodium 100mg/5ml if within safe limits how many ml will the nurse administer.

3. The nurse is caring for a child with a prescription for acetaminophen 68mg po every 4-6hours for pain or temperature greater than 100.5 F. The child weighs 10pounds the recommended dose of acetaminophen is 10-15mg/kg/dose every 4-6hrs

a. Is the prescribed amount a safe dose yes or  no explain

b. If with in safe limits how many ml will the nurse administer in 160mg/5ml

4. The nurse is caring for an infant weighing 2,600grams. The infant has been prescribed ceftazidime IV, the recommended dose for infants receiving ceftazidime is 30mg/kg every 12hr

a what is the safe amount per dose for this client

b. what is the safe daily dose for this client record your answers rounding to the nearest hundredth.

5. The nurse is caring for a client prescribed penicillin G 1.2 million units IM once STAT now the client is an adult weighing 110pounds using 5,000000units

a.How manyml of diluent will the nurse choose and why

b. How many ml will the nurse administer to the client

The Nurse is caring for a client with a prescription for 0.45% sodium chloride, 500ml IV to infuse over 8hrs. There is an infusion pump available how many ml/hr should the nurse administer to the clent record answer using a whole number

6. The client is caring for a cjild with a prescription for D5W, 1,800ml IV to infuse over 24hrs. There is an infusion pump available. How many ml/hr should the nurse administer to the client. record answer using a whole number

7. The nurse is caring for a client with a prescription for vancomycin 1 gram IL of 0.9% normal saline to infuse over 2hrs, there is aninfusion pump availablehow many ml/hr should the nurse administer to the client record answer using a whole number.

8. The nurse is caring for a client with a prescription for a bolus of 0.9% sodium chloride, 250ml IV to infuse over 30 minutes. There is an infusion pump available. How many ml/hr should the nurse administer to the client. Record answer using a whole number

9. The nurse is caring for a client wit a prescription for phenobarbital 150mg IV loading dose x 1 now. The pharmacy dispense Phenobabital 150ml of 0.9% sodium chloride with directions to infuse over 20 minutes. There is an infusion pump available. How many ml/hr should the nurse administer to the client using a whole number

10. The nurse is caring for a client with a prescription for Ringer lactate 9(RL) 1 L IV to infuse over 12hrs. Ther is a gravity set available with a drop factor 15gtt/ml. How many drops per minute gtt/min should the nurse administer to the client record answer using a whole number

11.The nurse is caring for a client with a prescription for D5 1/2 NS to infuse at 50ml/hr ther is a gravity set available with a drop factor of 20gtt/ml. How many drops per minute(gtt/min) should the nurse administer to the client record answer using a whole number

12. The nurse is caring for a clent with a prescription for D5NS to infuse at 100ml/hr. There is a gravity set available with a dropmfactormof 10gtt/ml. How many drops per minute (gtt/min) should the nurse administer to the client record answer using a whoel number

13.The nurse is caring for a client with a prescription for D5NSat 75ml/hr. There is a gravity set available with a drop factor of 60gtt/ml, how many drops per minute (gtt/min) should the nurse administer to the client. Record answer using a whole number
Answer #1

1. Ordered dose of Enoxaparin - 2mg/kg subcutaneous daily

Weight of the patient is 155lbs 8oz = 155×0.45kg = 70.3 kg

8oz to kg = 0.22kg

Total weight = 70.3 + 0.2kg = 70.5kg

Available Enoxaparin is 40mg/0.4ml

The dose required by the patient is 2× 70.5 = 141mg

Using unitary method :-

40mg of enoxaparin is available in 1ml

1 mg will be present in 1/40ml

Then 141mg will be present in 1/40× 141

= 3.5ml

Hence ,3.5ml of medicine has 141 mg of enoxaparin .

3.5 ml of medicine has to be given by the nurse to the patient.

2. Ordered dose of dicloxacillin sodium is 50mg po every 6hrly.

Weight of the child is 36lbs 4oz

36 lbs to kg = 36×0.45kg =16.2 kg

4oz is equal to 0.1 kg

Total weight - 16.3 kg

Recommended dosage - 12 to 13 mg /kg / day for children less than 40kg in every 6hours

So ,the dose for this patient is 16.3× 13= 211.9 mg /day which has to be divided into 4 doses to be given 6hourly which will be = 52 .9 mg can be given at one time . Hence the dose is safe

a) The prescribed dose is safe

b ) Available dicloxacillin sodium is 100mg /5ml

Using unitary method :-

100mg of dicloxacillin sodium is present in 5ml

Then 1mg will be present in 5/100ml

50mg will be present in 5/100×50

= 2.5ml has to be given by the nurse to the patient every 6 hourly .

3. Ordered dose of acetaminophen is 68mg po every 4-6hours for pain or temperature greater than 100.5 F

Weight of the child is 10lbs = 10×0.45= 4.5 kg

Recommended dose of acetaminophen is 10-15mg/kg/dose every 4-6hrs

Hence the dose for child is - 15× 4.5 = 67.5 mg every 4-6 hours

a) The dose is safe

b ) Available acetaminophen is 160mg/ 5ml

Using unitary method :-

160mg of acetaminophen is present in 5ml

1mg will be present in 5/160ml

Then 68mg will be present in 5/160× 68

= 2.1 ml

The nurse should give 2.1 ml of acetaminophen to the child every 4-6 hourly

4. Weight of the infant is 2600gm = 2.6kg

Recommended dose of ceftazidime is 30mg/kg every 12hourly

a) Safe dose for the infant is 30×2.6 = 78mg every 12hourly

b) Safe daily dose for this infant is 78×2 = 156mg /day